# How Fair is Your Diffusion Recommender Model?

### v1

@@ Abstract.
Diffusion-based recommender systems have recently proven to outperform traditional generative recommendation approaches, such as variational autoencoders and generative adversarial networks. Nevertheless, the machine learning literature has raised several concerns regarding the possibility that diffusion models, while learning the distribution of data samples, may inadvertently carry information bias and lead to unfair outcomes. In light of this aspect, and considering the relevance that fairness has held in recommendations over the last few decades, we conduct one of the first fairness investigations in the literature on DiffRec, a pioneer approach in diffusion-based recommendation. First, we propose an experimental setting involving DiffRec (and its variant L-DiffRec) along with nine state-of-the-art recommendation models, two popular recommendation datasets from the fairness-aware literature, and six metrics accounting for accuracy and consumer/provider fairness. Then, we perform a twofold analysis, one assessing models’ performance under accuracy and recommendation fairness separately, and the other identifying if and to what extent such metrics can strike a performance trade-off. Experimental results from both studies confirm the initial unfairness warnings but pave the way for how to address them in future research directions.
1. Introduction and motivation
@@
Research in recommender systems (RSs) has recently been shaped by the latest advances in deep learning, e.g., graph neural networks (Wu et al., 2023a), large language models (Wu et al., 2023b), and diffusion models (Wang et al., 2023). While the first two families have already received significant attention in the literature, diffusion-based RSs are still in early stages.
@@
As of today, diffusion models (Ho et al., 2020; Yang et al., 2024; Cao et al., 2024) have settled as the leading solution in generative AI (Goodfellow et al., 2020) and revolutionized the paradigm of generating new (but realistic) data with several applications in computer vision (Rombach et al., 2022; Ho et al., 2020), natural language processing (Austin et al., 2021; Li et al., 2022), graph data (Niu et al., 2020; Liu et al., 2023), chemistry, and bioinformatics (Xu et al., 2022; Hoogeboom et al., 2022). The core principle behind diffusion models lies in their twofold learning procedure, involving a forward process where samples from a real-data distribution are iteratively corrupted by adding noise, e.g., Gaussian noise, and a reverse process where the model is trained to reconstruct the original data by denoising the corrupted samples.
@@
As observed by Wang et al. (2023), the diffusion model framework is well-suited for recommendation settings, where it may effectively learn to denoise users’ implicit feedback containing noisy interactions, i.e., false positives. Their approach, called DiffRec, shows improved performance to traditional generative RSs, e.g., those based on variational autoencoders (Liang et al., 2018) and generative adversarial networks (Wang et al., 2017) by addressing their limited representational power (the former) or unstable training procedures (the latter). Similarly, other works have proposed to leverage diffusion models for specific recommendation tasks such as point-of-interests (Qin et al., 2024), sequential (Li et al., 2023; Ma et al., 2024), or knowledge graph-based recommendation (Jiang et al., 2024).
@@
Diffusion models have achieved widespread success in many domains and applications, but they are a relatively recent trend in machine learning, still facing several open challenges. For instance, some works have raised concerns regarding unfairness that could affect diffusion models. As observed in generative AI, bias is intrinsic in real-world data, and diffusion models might generate unfair outcomes while learning from such a biased data. In their seminal paper, Friedrich et al. (2023) discuss how stable diffusion methods are prone to generate synthetic images that lack diversity, e.g., firefighters are only associated with white males. Similarly, other works, such as Bianchi et al. (2023), discuss how textual prompts with trait annotations may produce stereotypical images, while Cho et al. (2023) observe bias in the generated images in terms of people color tones and occupations. Recent papers have proposed possible countermeasures, such as introducing fair guidance instructions during the data generation process (Friedrich et al., 2023), obfuscating the sensitive attributes (Choi et al., 2024), or adopting an ad-hoc fine-tuning (Shen et al., 2023).
@@
In light of the above, a natural question arises: “Can diffusion models be harmful to recommendation fairness likewise?”. The question is pertinent given the large relevance that bias and fairness have held in recommendations for decades (Anelli et al., 2022; Mansoury et al., 2022; Sun et al., 2020; Zhao et al., 2022b; Boratto et al., 2023b). Thus, we consider it is imperative to rigorously evaluate the performance of popular diffusion-based RSs, with a specific focus on potential unfairness issues, as there is currently no literature on this topic. Indeed, it is worth underlining that such recommendation models are still in their pioneer era. A thorough investigation of their performance can help identify and address observed pitfalls.
@@
To this end, in the current work, we select one of the pioneer approaches in diffusion-based recommendation, DiffRec (Wang et al., 2023) (and its variant L-DiffRec), to ultimately answer the following research question: “Is DiffRec a fair recommender system compared to the state-of-the-art?”. To conduct our investigation, we consider two datasets (with users’ attribute data) and nine additional state-of-the-art recommender systems. Our purpose is to benchmark their performance on six recommendation metrics accounting for accuracy, consumer (Boratto et al., 2022; Anelli et al., 2023; Vassøy and Langseth, 2024), and provider fairness (Boratto et al., 2024; Naghiaei et al., 2022; Mansoury et al., 2022). Concretely, we run a twofold performance analysis, where we first measure each performance metric separately, and then in a trade-off setting. We believe this twofold perspective can help identify performance patterns by encompassing all actors (consumers and providers) involved in the recommendation scenario (Anelli et al., 2023).
@@
Results depict a worrying scenario, outlining that fairness concerns raised in machine learning could also hold for the recommendation domain. Nonetheless, introducing additional components in the diffusion-based process, as for L-DiffRec, seems to mitigate the negative effects of unfairness; indeed, these findings pave the way to novel future directions in diffusion-based recommendation where fairness is preserved. Despite being a preliminary investigation, to the best of our knowledge, this is the first attempt in the literature to provide a rigorous and fairness-aware evaluation perspective on diffusion-based recommendation.
@@ 2. Background
This section provides the background notions regarding the two main topics of this paper, namely, diffusion-based recommendation (with a special focus on DiffRec (Wang et al., 2023)) and the evaluation metrics to assess fairness (consumer and provider sides) in recommendation.
@@ 2.1. Diffusion recommender model
Similarly to diffusion models, DiffRec (Wang et al., 2023) proposes to treat recommendation as a denoising process of historical user-item interactions, which may be extremely noisy, i.e., contain false positives. In detail, the recommendation model is trained to first corrupt users’ interactions through noise (forward process) and then reconstruct them via an ad-hoc denoising procedure (reverse process). Differently from traditional diffusion models employed on images, users’ historical interactions are not corrupted into pure noise, but the ratio of added noise decreases consistently.
@@
Let and be the sets of users and items in the system, respectively, with as the vector of user-item interactions for user . Each vector entry denotes user interacted with item , otherwise . Then, the forward process, in a Markov chain process, corrupts the user interaction vector into steps by gradually adding Gaussian noise. The formulation for any step is: , where (with a notation abuse) we indicate and set , , at the initial step, while stands for the Gaussian distribution, controls the magnitude of the Gaussian noise added at each step , and is the standard deviation of the distribution. Conversely, the reverse process involves training any machine learning model, e.g., a neural network, to approximate the mean and standard deviation of the data distribution and recover the original user interaction vector. Again, for any step , we have: , where and are the learned mean and standard deviation through the machine learning model having parameters .
@@
Furthermore, to address large-scale recommendation (a pretty common scenario in real-world applications), Wang et al. (2023) also propose L-DiffRec, a lighter-weight version of DiffRec. This variant first clusters pre-trained item embeddings, e.g., from LightGCN, into clusters. Afterward, the initial user interaction vector is split into sub-vectors according to the corresponding -th item cluster. Then, a variational autoencoder is trained to compress each of user interaction sub-vectors , so that the diffusion, i.e., forward/reverse processes, occurs in a low-dimensional latent space, ideally in parallel computation across all splits. Finally, the original user interaction vector is reconstructed through the decoding process of the previous variational autoencoder.
@@
Note that in Wang et al. (2023), the authors eventually propose a second variant to DiffRec, called T-DiffRec, which can encode the temporal information during the training. However, T-DiffRec is specialized for sequential recommendation, which would require a dedicated study to compare its performance with other analogous models. To this end, we exclude T-DiffRec from our experiments, leaving its adoption to future extensions of our study.
@@
Our analysis aims to offer a preliminary outline on the fairness performance of diffusion-based recommender systems. Although other works have recently contributed with novel diffusion-based solutions (Qin et al., 2024; Li et al., 2023; Ma et al., 2024; Jiang et al., 2024), DiffRec could be rightly considered as the pioneer approach in this domain for chronological reasons, In light of this, we selected only this model and plan to extend our investigation to other diffusion models in forthcoming works.
@@ 2.2. Fairness evaluation in recommendation
Recent studies focused on assessing fairness under the notion of demographic parity (Boratto et al., 2022; Naghiaei et al., 2022; Vassøy and Langseth, 2024). On the one hand (Boratto et al., 2023a; Naghiaei et al., 2022), we operationalize demographic parity on the consumer side as the absolute difference in recommendation utility between users’ groups. We select the respective fairness counterparts of nDCG and Recall, namely nDCG and Recall. On the other hand, provider fairness is estimated as the absolute difference in exposure between items’ groups (Exp) (Naghiaei et al., 2022; Boratto et al., 2024) and the average percentage of long-tail items in the top- lists (APLT) (Anelli et al., 2023; Mansoury et al., 2022). Lower values of these metrics reflect fairer recommendations, except for APLT. For page constraints, we do not dive deep into their formulations. The interested reader may refer to the above-cited papers for a detailed presentation.
@@ 3. Experimental study
In this section, we present our proposed experimental study aimed at assessing fairness in diffusion-based recommendation. First, we describe the baselines and datasets adopted for the analysis. Then, we provide reproducibility details for the conducted experiments.
@@ 3.1. Baselines
We select nine state-of-the-art recommendation models along with DiffRec and L-DiffRec: BPRMF (Rendle et al., 2009), ItemkNN (Sarwar et al., 2001), NeuMF (He et al., 2017), LightGCN (He et al., 2020), UltraGCN (Mao et al., 2021), XSimGCL (Yu et al., 2024), EASE (Steck, 2019), MultiVAE (Liang et al., 2018), and RecVAE (Shenbin et al., 2020). As reported in Table 1, our selection depends on the following criteria: (i) the baselines are well-established solutions, accepted at top-tier conferences and journals; (ii) they cover a diverse set of recommendation families, such as graph neural networks and generative models; (iii) many approaches have been already analyzed under recommendation fairness, highlighting an existing literature gap in diffusion-based recommendation.
@@ 3.2. Datasets
We assess the performance of the selected baselines on MovieLens-1M (ML1M) (Harper and Konstan, 2016) and Foursquare Tokyo (FTKY) (Yang et al., 2016), which are frequently adopted in fairness-aware literature in recommendation (Vassøy and Langseth, 2024; Boratto et al., 2023a; Liu et al., 2021). While ML1M is a popular recommendation dataset that collects movies and users’ preferences on them, FTKY involves interactions represented by users’ check-ins at points of interest in Tokyo. In both datasets, users’ sensitive information comes in a binary gender label. Thus, fairness analyses on the consumer side adopt the gender attribute included in both datasets, whereas provider fairness is assessed on a binary partition of the items’ set, namely short-head (popular) and long-tail (niche) items. Following (Boratto et al., 2024; Naghiaei et al., 2022; Mansoury et al., 2022), we select the 20% most popular items as short-head and the rest as long-tail. Dataset statistics are reported in Table 2.
3.3. Reproducibility
Datasets are split into train, validation, and test sets (70%/10%/20%) following the temporal hold-out method. Then, for the models training and evaluation, we rely on the framework RecBole (Zhao et al., 2022a) for training and evaluating all the baselines and, for the two diffusion-based recommender systems, we adopt their original code implementations111https://github.com/YiyanXu/DiffRec.. All recommendation baselines are trained with a grid-search exploration of the hyper-parameters by following the original papers and selecting the hyper-parameter configuration providing the highest value of the Recall@20 on the validation set.
| Models | Year | Venue | Analyzed fairness in |
@@ 3.3. Reproducibility
| ItemkNN (Sarwar et al., 2001) | 2001 | WWW | (Anelli et al., 2022; Boratto et al., 2023a, b, 2022) |
| NeuMF (He et al., 2017) | 2017 | WWW | (Anelli et al., 2022; Boratto et al., 2023a, b; Mansoury et al., 2022; Sun et al., 2020; Boratto et al., 2022) |
| LightGCN (He et al., 2020) | 2020 | SIGIR | (Anelli et al., 2023; Boratto et al., 2023a; Zhao et al., 2022b) |
| UltraGCN (Mao et al., 2021) | 2021 | CIKM | (Anelli et al., 2023) |
| XSimGCL (Yu et al., 2024) | 2024 | TKDE | ✗ |

### v2

@@ Abstract.
Diffusion-based learning has settled as a rising paradigm in generative recommendation, outperforming traditional approaches built upon variational autoencoders and generative adversarial networks. Despite their effectiveness, concerns have been raised that diffusion models - widely adopted in other machine-learning domains - could potentially lead to unfair outcomes, since they are trained to recover data distributions that often encode inherent biases. Motivated by the related literature, and acknowledging the extensive discussion around bias and fairness aspects in recommendation, we propose, to the best of our knowledge, the first empirical study of fairness for DiffRec, chronologically the pioneer technique in diffusion-based recommendation. Our empirical study involves DiffRec and its variant L-DiffRec, tested against nine recommender systems on two benchmarking datasets to assess recommendation utility and fairness from both consumer and provider perspectives. Specifically, we first evaluate the utility and fairness dimensions separately and, then, within a multi-criteria setting to investigate whether, and to what extent, these approaches can achieve a trade-off between the two. While showing worrying trends in alignment with the more general machine-learning literature on diffusion models, our results also indicate promising directions to address the unfairness issue in future work. The source code is available at https://github.com/danielemalitesta/FairDiffRec.
1. Introduction
@@
Research in recommender systems (RSs) has recently been shaped by the latest advances in deep learning, e.g., graph neural networks (Wu et al., 2023), language models (Balloccu et al., 2023; Wu et al., 2024), and diffusion models (Wang et al., 2023). While the first two families have already received significant attention in the literature, diffusion-based RSs are still in early stages.
@@
As of today, diffusion models (Ho et al., 2020; Yang et al., 2024) have settled as the leading solution in generative artificial intelligence (AI) (Goodfellow et al., 2020) and revolutionized the paradigm of generating new (but realistic) data with several applications, such as computer vision (Rombach et al., 2022; Ho et al., 2020), natural language processing (Austin et al., 2021; Li et al., 2022), and graph modeling (Niu et al., 2020; Liu et al., 2023), and domains, such as chemistry and bioinformatics (Xu et al., 2022; Hoogeboom et al., 2022). The core principle behind diffusion models lies in their twofold learning procedure, involving a forward process where samples from a real-data distribution are iteratively corrupted by adding noise, e.g., Gaussian noise, and a reverse process where the model is trained to reconstruct the original data by denoising the corrupted samples.
@@
As observed by Wang et al. (2023), the diffusion model framework is well-suited for recommendation settings, as it can learn to denoise users’ implicit feedback containing noisy interactions, i.e., false positives. This approach, called DiffRec, shows improved performance with respect to traditional generative RSs, e.g., those based on variational autoencoders (VAEs) (Liang et al., 2018) and generative adversarial networks (GANs) (Wang et al., 2017), by addressing the limited representational power of the former and the unstable training procedures of the latter. Similarly, other works proposed to leverage diffusion models for specific recommendation tasks such as point-of-interests (Qin et al., 2024), sequential (Li et al., 2024b; Ma et al., 2024), or knowledge graph-based recommendation (Jiang et al., 2024).
@@
Diffusion models have achieved widespread success in many applications and domains; however, as a relatively recent development in machine learning (ML), they continue to face several open challenges. For instance, some works have raised concerns regarding the unfair outcomes diffusion models could lead to. As observed in generative AI, bias is intrinsic in real-world data, and diffusion models might amplify this signal while learning the underlying biased data distribution. In their seminal paper, Friedrich et al. (2023) discuss how stable diffusion methods are prone to generate synthetic images that lack diversity, e.g., firefighters are only associated with white males. Other works, such as Bianchi et al. (2023), discuss how textual prompts with trait annotations may produce stereotypical images. Similarly, Cho et al. (2023) observe bias in the generated images in terms of people color tones and occupations. Interestingly, the same unfairness issues have also been observed in graph diffusion models (Vignac et al., 2023; Chen et al., 2023), where the generation of both the graph topology and the node features (Li et al., 2024a) further amplifies the natural bias already existing in the graph structure (Kose and Shen, 2024). Recent papers have proposed possible countermeasures, such as introducing fairness-guided instructions during the data generation process (Friedrich et al., 2023), obfuscating the sensitive attributes (Choi et al., 2024), or adopting an ad-hoc fine-tuning (Shen et al., 2024; Kose and Shen, 2024).
@@
In this light, a question arises: “To what extent do diffusion models exacerbate or replicate unfairness patterns within recommender systems?” This question is pertinent given the large relevance that bias and fairness have held in recommendations for decades (Anelli et al., 2022; Mansoury et al., 2022; Sun et al., 2020; Zhao et al., 2022b; Boratto et al., 2023b). Thus, we consider it imperative to evaluate the performance of popular diffusion-based RSs rigorously, with a specific focus on potential unfairness issues, as there is currently no literature on this topic, despite a well-documented history of bias and unfairness affecting other ML-based RSs. As diffusion-based recommender models are still in their pioneer era, a thorough investigation of their performance can help identify and address observed pitfalls.
@@
To this end, in this work, we select one of the pioneer approaches in diffusion-based recommendation, DiffRec (Wang et al., 2023) (and its variant L-DiffRec), to answer the following research question: “Is DiffRec a fair recommender system compared to the state-of-the-art?” To conduct our investigation, we consider two datasets (with users’ attribute data and binary gender label as the sensitive attribute) and nine additional state-of-the-art recommenders. Our purpose is to benchmark their performance on six recommendation metrics accounting for utility, consumer fairness (Boratto et al., 2022; Anelli et al., 2023; Vassøy and Langseth, 2024), and provider fairness (Boratto et al., 2024b; Naghiaei et al., 2022; Mansoury et al., 2022). Concretely, we run a twofold analysis, where we first monitor each dimension separately, and then in a multi-criteria setting. This twofold perspective can help identify patterns by considering all actors in the recommendation scenario (Anelli et al., 2023).
@@
Our results depict a worrying scenario, confirming that unfairness concerns already observed for other ML tasks do indeed extend to the recommendation task, when leveraging diffusion-based approaches. Nonetheless, introducing additional components in the diffusion-based process, as for L-DiffRec, seems to mitigate the observed unfairness; indeed, these findings pave the way to novel future directions where fairness is preserved in diffusion-based recommendation. Despite its preliminary nature, to the best of our knowledge, this is the first attempt to provide a rigorous, fairness-aware evaluation of diffusion-based recommendation methods.
@@ 2. Background
This section provides the necessary notions regarding diffusion-based recommendation (with a special focus on DiffRec (Wang et al., 2023)), and consumer fairness and provider fairness in recommendation.
@@ 2.1. Diffusion Recommender Model
Building upon diffusion models, DiffRec (Wang et al., 2023) proposes to treat the recommendation task as a denoising process of historical user-item interactions, which may be overly noisy (i.e., contain false positives). In detail, the model is trained to first corrupt users’ interactions through noise (forward) and then reconstruct them via an ad-hoc denoising procedure (reverse). Unlike traditional diffusion models for images, users’ historical interactions are not directly corrupted into pure noise, but the ratio of added noise decreases consistently.
@@
Let and be the sets of users and items in the system, respectively, with as the vector of user-item interactions for user . Each vector entry denotes user interacted with item , otherwise . Then, the forward process, in a Markov chain process, corrupts the user interaction vector into steps by gradually adding Gaussian noise. The formulation for any step is: , where (with a notation abuse) we indicate , , at the initial step, stands for the Gaussian distribution, controls the magnitude of the Gaussian noise added at each step , and is the standard deviation of the distribution. Conversely, the reverse process involves training any ML model, e.g., a neural network, to approximate the mean and standard deviation of the data distribution and recover the original user interaction vector. Again, for any step , we have: , where and are the learned mean and standard deviation through the ML model parametrized by .
@@
Furthermore, to address large-scale recommendation (a pretty common scenario in real-world applications), Wang et al. (2023) also propose L-DiffRec, a lighter-weight version of DiffRec. This variant first clusters pre-trained item embeddings, e.g., from LightGCN, into clusters. Afterward, the initial user interaction vector is split into sub-vectors according to the corresponding -th item cluster. Then, a VAE is trained to compress each user interaction sub-vector , so that the diffusion, i.e., forward/reverse processes, occurs in a low-dimensional latent space, ideally in parallel computation across all splits. Finally, the original user interaction vector is reconstructed through the decoding process of the previous VAE.
@@
Note that in Wang et al. (2023), the authors eventually propose a second variant to DiffRec, called T-DiffRec, which can encode the temporal information during the training. However, T-DiffRec is specialized for sequential recommendation, which would require a dedicated study to compare its performance with other analogous models. To this end, we exclude T-DiffRec from our experiments, leaving its for future extensions of our study.
@@
Our analysis aims to offer a preliminary outline on the fairness performance of diffusion-based RSs. Although other works have recently contributed with novel diffusion-based solutions (Qin et al., 2024; Li et al., 2024b; Ma et al., 2024; Jiang et al., 2024), DiffRec could be rightly considered as the pioneer approach in this domain for chronological reasons. In light of this, we plan to extend our analysis to other diffusion models in future works.
@@ 2.2. Fairness Evaluation in Recommendation
Recent studies, both in RS (Boratto et al., 2022; Naghiaei et al., 2022; Vassøy and Langseth, 2024; Boratto et al., 2023a; Medda et al., 2025) and relatedly user modeling (Purificato et al., 2025, 2022, 2024), focused on assessing algorithmic fairness under the notion of demographic parity (Dwork et al., 2012; Feldman et al., 2015). From a consumer perspective (Boratto et al., 2024a; Naghiaei et al., 2022), we operationalize this notion as the absolute difference in recommendation utility between users’ groups. To this end, we select the respective fairness counterparts of nDCG and Recall, namely nDCG and Recall. On the other hand, provider fairness is estimated as the absolute difference in exposure (i.e., how evenly items or groups of items are exposed to users or groups of users) between items’ groups (Exp) (Naghiaei et al., 2022; Boratto et al., 2024b) and as the average percentage of long-tail items in the top- lists (APLT) (Anelli et al., 2023; Mansoury et al., 2022). For all metrics except APLT, values closer to zero indicate fairer recommendations. Due to space constraints, we do not delve into the detailed formulations. Interested readers are referred to the cited works for a comprehensive presentation.
@@ 3. Experimental Study
In this section, we present our experimental study aimed at assessing recommendation utility and fairness in diffusion-based methods. First, we describe the baselines and datasets adopted for the analysis. Then, we provide reproducibility details for the experiments.
@@ 3.1. Baselines
We select nine state-of-the-art recommendation models along with DiffRec and L-DiffRec: BPRMF (Rendle et al., 2009), ItemkNN (Sarwar et al., 2001), NeuMF (He et al., 2017), LightGCN (He et al., 2020), UltraGCN (Mao et al., 2021), XSimGCL (Yu et al., 2024), EASE (Steck, 2019), MultiVAE (Liang et al., 2018), and RecVAE (Shenbin et al., 2020). As reported in Table 1, our selection was guided by the following observations: (i) the baselines are well-established solutions, accepted at top-tier conferences and journals; (ii) they cover a diverse set of recommendation families, such as graph neural networks and generative models; (iii) many of these approaches have been already analyzed from a recommendation fairness perspective.
@@ 3.2. Datasets
We assess the performance of the selected baselines on MovieLens-1M (ML1M) (Harper and Konstan, 2016) and Foursquare Tokyo (FTKY) (Yang et al., 2016), which are frequently adopted in fairness-aware literature in recommendation (Vassøy and Langseth, 2024; Boratto et al., 2024a; Liu et al., 2021). While ML1M is a popular recommendation dataset that collects movies and users’ preferences on them, FTKY includes users’ check-ins at points of interest in Tokyo. In both datasets, users’ sensitive information comes in a binary gender label. Thus, fairness analyses on the consumer side adopt the gender attribute included in both datasets (where female is the protected group), whereas provider fairness is assessed on a binary partition of the items’ set, namely short-head (popular) and long-tail (niche) items (where niche items are the protected group). Following a standard procedure in the literature, as in (Boratto et al., 2024b; Naghiaei et al., 2022; Mansoury et al., 2022), we select the 20% most popular items as short-head and the rest as long-tail. Dataset statistics are reported in Table 2.
3.3. Reproducibility Details
Datasets are split into train, validation, and test sets (70%/10%/20%) following the temporal hold-out method adopted in (Wang et al., 2023; Boratto et al., 2024a). We use the RecBole framework (Zhao et al., 2022a) to train and evaluate all baseline models. For the two diffusion-based recommender systems, we adopt their original code implementations111https://github.com/YiyanXu/DiffRec. Accessed August 10, 2025.. Baseline models are trained using a grid search over hyperparameters, guided by the settings in the original papers. The configuration yielding the highest Recall@20 on the validation set is selected for each model.
| Models | Year | Venue | Fairness analyzed in |
| ItemkNN (Sarwar et al., 2001) | 2001 | WWW | (Anelli et al., 2022; Boratto et al., 2023b, 2022) |
@@ 3.3. Reproducibility Details
| NeuMF (He et al., 2017) | 2017 | WWW | (Anelli et al., 2022; Boratto et al., 2023b; Mansoury et al., 2022; Sun et al., 2020; Boratto et al., 2022) |
@@
| LightGCN (He et al., 2020) | 2020 | SIGIR | (Anelli et al., 2023; Boratto et al., 2024a; Zhao et al., 2022b) |
| UltraGCN (Mao et al., 2021) | 2021 | CIKM | (Anelli et al., 2023; Boratto et al., 2024a) |